OpenStudy (scarlettfarra2000):

Will Medal Simplify -24 divided by 1/6

OpenStudy (astrophysics):

Take second fraction flip and multiply

OpenStudy (anonymous):

dividing by one sixths is the same as multiplying by six

OpenStudy (scarlettfarra2000):

So the answer would be -4?

OpenStudy (astrophysics):

What's -24 x 6?

OpenStudy (scarlettfarra2000):

-144

OpenStudy (astrophysics):

Done

OpenStudy (scarlettfarra2000):

Thanks so much

OpenStudy (astrophysics):

\[\frac{ a }{ b } \div \frac{ c }{ d } \implies \frac{ a }{ b } \times \frac{ d }{ c }\]